OpenStudy (anonymous):

A rectangular patio is 9 ft by 6 ft. When the length and width are increased by the same amount, the area becomes 88 sq ft. Ginger is using the zero product property to solve the equation (6 + x)(9 + x) = 88. What do her solutions represent? a.one possible length and one possible width b.two possible amounts for the widths c.two possible amounts for the lengths d.two possible amounts by which the dimensions were increased

OpenStudy (shinalcantara):

(6 + x)(9 + x) = 88 \[x^2 + 15x + 54 = 88\] Since this is in quadratic form, then you'll get two values of x

OpenStudy (shinalcantara):

\[x^2 + 15x +54 = 88\] \[x^2 + 15x +54 -88 = 0\] \[x^2 + 15x -34 = 0\] \[(x-2)(x+17)=0\] \[(x-2)=0 ; (x+17)=0\] \[x=2;x=-17\]

OpenStudy (shinalcantara):

take the positive value since you're getting the length that was added and you can't have a negative answer for that.

OpenStudy (shinalcantara):

x represents the length increased in the dimension therefore D is the answer

OpenStudy (anonymous):

thank you.

OpenStudy (shinalcantara):

yw :)

OpenStudy (anonymous):

D